OpenStudy (anonymous):

Find the range of y=(1/x+4)-1

OpenStudy (coconutjj):

First find the range of 1/x+4

OpenStudy (anonymous):

okay

OpenStudy (coconutjj):

the numerator is 1, what value can the fraction not equal to

OpenStudy (anonymous):

uh i'm not sure

OpenStudy (coconutjj):

well if its 0/x+4 what is our solution?

OpenStudy (anonymous):

-4/x? I really don't know. This is why I dropped out of math. but I have to finish my crcedit recovery for it

OpenStudy (anonymous):

range is the value of y

OpenStudy (anonymous):

so what value of x that would make the value y infinite

OpenStudy (anonymous):

if there isn't any then R is answer if there is the R except that value

OpenStudy (anonymous):

1/(x+4)-1 this equation could be equal infinite if the denominator is equal to zero so x+4=0 =>x=-4

OpenStudy (xapproachesinfinity):

hmm first find the domain since range depends on what the domain is \[y=\frac{1}{x+4}-1\] the domain is all x except -4 \[( -\infty, -4)\cup (-4, \infty)\] now at \[x=4 ~we~have~ y=\pm\infty \] at \[x=\pm\infty \Longrightarrow y=-1\]

OpenStudy (xapproachesinfinity):

see what you can conclude from this!

OpenStudy (xapproachesinfinity):

really people should be asked range if there is no graph because finding the range involve calculus sometimes for some function or the use of inverse functions if they exist! the goal is to see where the domain maps to

OpenStudy (xapproachesinfinity):

shouldn't be*

OpenStudy (xapproachesinfinity):

if you graph this thing you should be able to see the range clearly

OpenStudy (xapproachesinfinity):

see attached! http://prntscr.com/6luml1